OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). a. y = -2x - 4 b. y = -2x + 18 c. y = -2x - 8 d. y = -2x – 12

OpenStudy (ciarán95):

Given a line passing through a point with coordinates (x1, y1) and having a slope of m, the equation of that line can be given by: \[y - y _{1} = m(x - x _{1})\] So, by substituting in our values for x1, y1 and m into the above and rearranging to get y on its own, you should have the answer. Hope that helps! :)

OpenStudy (anonymous):

Can you show me how lol

OpenStudy (ciarán95):

For our line, we have a point (2, -8) that it passes through and a slope of -2. We now need to label our x1, y1 and m, as I talked about in my previous post. So, x1 = 2 x2 = -8 m = -2 Now, replace the x1, y1 and m in our general equation for the line with 2, -8 and -2 respectively. This will give us: \[y -(-8) = -2(x -2)\] or \[y + 8 = -2(x - 2)\] Then, it's a case of multiplying out the bracket on the right-hand side of the equation and isolating the 'y' term on its own on the left-hand side so that your equation looks like one of the options listed in the question.

OpenStudy (anonymous):

y = -2x -4 a.